A voltaic cell consists of a Zn/Zn2+ anode and a Ni/Ni2+ cathode at 25 ∘C. The initial concentrations of Ni2+ and Zn2+ are 1.700 M and 0.120 M , respectively. The volume of half-cells is the same. What is the concentration of Ni2+ when the cell potential falls to 0.457 V ? Enter your answer to 4 decimal places and in units of mM.
Answer #1

A node :- cathode :- zn Ni2+ + 2e ent the Nics) Zn (s) + Ni² Ni cay, zn cap, + Nics) Ell = Ered - EOK = -0.25-(-0.763) = 0.